OpenStudy (ineedhelp64):

Algebra 2 Question- Using complete sentences, explain how to find the minimum value for each function and determine which function has the smallest minimum y-value. f(x) = 3x2 + 12x + 16 g(x) = 2 * sin(x - pi)

OpenStudy (welshfella):

Have you done any calculus - finding derivatives?

OpenStudy (ineedhelp64):

no

rebeccaxhawaii (rebeccaxhawaii):

This is trig?

OpenStudy (ineedhelp64):

Yeah

rebeccaxhawaii (rebeccaxhawaii):

Thought so.

OpenStudy (ineedhelp64):

im really bad at trig

OpenStudy (welshfella):

you can find the minimum value of the first function by completing the square

rebeccaxhawaii (rebeccaxhawaii):

I'm in the same boat as you, buddy. I'm taking it rn. ;-; You probably know more about it than me. Good luck to you!

OpenStudy (ineedhelp64):

Can you please explain to me how to do that?

OpenStudy (welshfella):

OK You have to convert the function to the vertex form a(x - b)^2 + c where c will be the minimum value f(x) = 3x^2 + 12x + 16 divide first 2 terms by 3 = 3(x^2 + 4x) + 16 now x^2 + 4x = (x + 2)^2 - 4 so we have 3[(x + 2) ^2 - 4] + 16 = 3(x + 2)^2 + 16 - 12 = 3(x + 2)^2 + 6 so the minimum value is 4.

OpenStudy (welshfella):

* penultimate line should read + 4 NOT + 6

OpenStudy (welshfella):

so minimum value of f(x) is 4 and you can read the minimum value of g(x) from the graph

OpenStudy (ineedhelp64):

Thanks that really did help me out a lot. I appreciate it

OpenStudy (welshfella):

To find the minimum value of g(x) without the graph you use the fact that the minimum value of the sine function is -1 . So since you have 2 sin (x - pi) the minimum value is 2*-1.

OpenStudy (welshfella):

yw